# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 4.1.1. SLM Applications in Question-Answering
    Retrieval-Augmented Generation (RAG): Incorporating External Knowledge for Domain-Specific QA Retrieval-augmented generation (RAG) addresses out-of-domain (OOD) questions by integrating external knowledge during inference, allowing models to access information beyond their pre-trained parameters. By retrieving relevant documents in real time, RAG enables small language models to provide accurate answers on specialized topics. In the telecommunications domain, Gichamba et al. (2024) use ColBERT as a dense retrieval system to fetch documents from technical datasets. By encoding queries and documents separately, ColBERT computes relevance scores, helping small models like Phi-2 and Falcon-7B retrieve precise technical information to answer complex telecom-related queries. Hartill et al. (2023b) propose an approach combining multi-task pre-training with dense retrieval for compositional questions. The model decomposes complex, multi-hop queries and retrieves evidence from external sources, showing improved performance on benchmarks like CommonsenseQA (Talmor et al., 2019) and ARC-DA (Clark et al., 2018). EffiChainQA (Roh et al., 2024) introduces a chain-of-reasoning framework that enhances reasoning in open-domain QA through retrieval and question decomposition. By breaking down queries and retrieving documents using a retriever fine-tuned on HotpotQA (Yang et al., 2018), EffiChainQA achieves performance comparable to state-of-the-art models on datasets like HotpotQA, demonstrating the efficacy of SLMs in managing complex tasks with lower computational costs.
@@
    Self-Adaptive Techniques: Enhancing Model Adaptability with Self-Generated Pseudo Labels Despite LMs’ capacity to store extensive general knowledge applicable across various tasks, they often exhibit suboptimal performance when transferring and adapting this knowledge to specific downstream tasks. Fine-tuning, while effective, can be impractical in realistic scenarios where labeled datasets are scarce. To overcome this, self-adaptive techniques employ self-generated pseudo labels to activate specific aspects of the target tasks, thereby enhancing model adaptability (Veksler, 2023; Shu et al., 2022). Jeong et al. (2023) introduce Test-time Self-Adaptive Small LMs (T-SAS), a framework for SLMs that first stochastically generates multiple answers for an unlabeled question. The most plausible answer is then selected via majority voting to enhance pseudo-label accuracy, and samples with low agreement are filtered out. Evaluated on QA datasets like Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016), T-SAS shows improved model adaptability. For example, on the SQuAD dataset, Flan-T5 base (250M) (Chung et al., 2024) attains a 63.02 EM score, nearly matching the 69.94 score achieved by the fine-tuned Flan-T5 base with a training set.
  3. (3)
    Chain-of-Thought (CoT) Prompting: Generating Intermediate Reasoning Steps Chain-of-thought (CoT) prompting enhances small language models’ ability to handle out-of-domain (OOD) questions by encouraging them to generate intermediate reasoning steps. By articulating step-by-step thinking, CoT enables models to navigate unfamiliar topics more effectively and arrive at accurate answers. Rationale Ranking (Hartill et al., 2023a) propose to use an SLM that generates intermediate reasoning steps, or rationales, to provide answers grounded in evidence. By generating and ranking these rationales, the SLM supports its answers with a clear chain of reasoning, reducing the risk of hallucination when faced with unfamiliar or out-of-domain questions. Rationale ranking further enhances this process by scoring the generated rationales based on relevance and truthfulness. This ranking allows the model to prioritize the best rationale and improve accuracy. Together, these techniques ensure the model delivers reliable answers by systematically selecting the most appropriate reasoning pathway, even when encountering unfamiliar queries.
Comparison between LLMs and SLMs for QA. When comparing LLMs like GPT-4 (Achiam et al., 2023) or BLOOM-175B (Le Scao et al., 2023) with fine-tuned SLMs in QA tasks, the benefits of SLMs are clear. LLMs, while versatile across multiple domains due to extensive pre-training, are computationally demanding, making them less ideal for resource-limited settings. SLMs, however, when fine-tuned for specific domains, often match or exceed the performance of larger models within those specialties. The trade-off is between scale and specialization: LLMs handle diverse domains but may need additional techniques like Retrieval-Augmented Generation (RAG) for domain-specific queries. In contrast, domain-specific SLMs, though less flexible, provide higher accuracy and more relevant responses, making them ideal for edge deployments where computational resources are scarce but domain precision is crucial.
@@ 4.1.2. SLM Applications in Coding
The adoption of SLMs for coding offers an alternative to LLMs due to their lower computational needs and potential for domain-specific tuning. Despite LLMs’ proficiency in code generation and programming support, SLMs are advantageous for their faster inference, reduced operational costs, and suitability for real-time environments where rapid responses are crucial. Representative works are discussed next. The Phi series (Javaheripi et al., 2023; Abdin et al., 2024; Li et al., 2023b) showcase SLMs’ evolution in coding tasks. For instance, Phi-1 (Gunasekar et al., 2023), a Transformer with 1.3 billion parameters, specializes in basic Python coding and achieves notable scores in benchmarks such as HumanEval (Gunasekar et al., 2023), which includes 164 programming problems. Subsequent models, Phi-1.5 and Phi-2, have enhanced these capabilities, while Phi-3 demonstrated SLMs’ potential to rival larger models (Abdin et al., 2024). The latest model, Phi-3.5-mini, with 3.8 billion parameters, excels in long context tasks using advanced fine-tuning and optimization techniques, performing comparably to larger models such as Llama-3.1-8B-instruct (Dubey et al., 2024) and surpassing smaller ones like Gemma-2 (Team et al., 2024b).
@@
Comparison between SLMs and LLMs on Coding. Table 6 provides a comparative analysis of SLMs and LLMs on coding benchmarks HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). Insights include: (i) Small SLMs (1.3B - 3.8B Parameters) like Phi-3.5-mini (Trufinescu, 2024) achieve high scores, demonstrating the efficacy of small models. Mid-sized SLMs (6.7B - 9B Parameters), such as DeepSeek-Coder 6.7B (Guo et al., 2024b) and Llama 3.1 8B (Dubey et al., 2024), show improved performance, indicating that larger model sizes and enhanced training contribute to better accuracy. Large models (33B and above) like Llama 3.1 405B (Dubey et al., 2024), GPT-4o (OpenAI, 2024), and Claude 3.5 Sonnet (Anthropic, 2024) excel, supporting the idea that bigger models generalize better across diverse coding tasks; (ii) There’s a notable trade-off between computational efficiency and performance, with larger models requiring more resources, impacting their practical deployment in constrained environments; (iii) Specialized training and fine-tuning, as used in models like DeepSeek-Coder (Guo et al., 2024b), are crucial for excelling in coding tasks, though such models may not handle complex requests as effectively, highlighting the versatility of general SLMs for broader applications.
@@ 4.1.3. SLM Applications in Recommender Systems
Recommender systems are essential in various online services, helping to manage information overload and meet users’ needs. SLMs enhance recommendation systems by (1) addressing the cold start problem; (2) reducing popularity bias; (3) improving long-term planning; (4) serving as personalized recommenders; and (5) acting as content encoders. These applications show the versatility and effectiveness of SLMs in boosting performance and personalization in recommendation. Next, we introduce the details.
@@
SLM for System Cold Start Problem. Traditional recommendation systems, which utilize historical user-item interactions like clicks, purchases, and ratings to learn representations and match items to users, fail in scenarios lacking any user-item interactions, known as the cold-start recommendation problem, often occurring in start-up businesses (Rashid et al., 2008). Although LLMs address this with in-context learning, their slow and costly inference restricts real-time use. Thus, PromptRec (Wu et al., 2024d) explores using SLMs as in-context recommenders for recommendation system cold-start problems. However, SLMs often struggle without emergent context-learning abilities. To overcome this, SLMs are enhanced by pre-training on relevant corpora measured by mutual information between documents and user-item interactions, using a C4 corpus subset (Raffel et al., 2020), and by developing training prompts for different domains, enhancing cold-start performance. Results show that enhanced SLMs like BERT-mini (Devlin et al., 2019), with 11.3M parameters, achieve BERT-large’s performance in cold-start scenarios, with only 17% of BERT-large’s inference time. Similarly, many studies have addressed the cold-start problem by leveraging BERT (Zhang et al., 2019; Noorian, 2024; Zhuang and Kim, 2021; Heidari et al., 2022). For example, ADLRS (Heidari et al., 2022) employs BERT to convert web-crawled item profiles into vectors that highlight key aspects, aiding recommender systems in acquiring essential initial information.
Recent studies have employed LLMs to address the cold-start problem. For instance, Sanner et al. (2023) uses PaLM 62B (Chowdhery et al., 2023) with prompts to infer preferences for cold users based on their textual profiles, LLM Interaction Simulator (LLM-InS) (Huang et al., 2024c) uses Llama 2 (size unspecified) to simulate interactions and generate user interaction sets for cold items, enabling simultaneous training of cold and warm items to bridge their embedding gap, and Wang et al. (2024e) applies PaLM to determine user preferences for cold-start items from textual descriptions of users’ historical behaviors and new item descriptions. Despite the limited exploration of SLMs in the LLM era for recommendation cold-start problems, researching them is worthwhile due to their enhanced in-context learning capabilities and small size.
SLM for Mitigating Popularity Bias. Popularity bias in recommender systems, characterized by a discrepancy between item popularity in training datasets and the open world, often arises from training on closed-loop datasets that contain narrow and fixed information. Recent LLMs utilize their expansive open-world knowledge to enhance reasoning about user-item interactions (Lin et al., 2024b; Liu et al., 2024a), thereby mitigating this bias by enriching the inputs to recommenders with more comprehensive item information. However, the substantial resource demands of LLMs limit their practical application. To address this challenge, the Step-by-step Knowledge Distillation Framework for Recommendation (SLIM) (Wang et al., 2024f) distills the reasoning capabilities of LLMs into smaller language models (SLMs), retaining only 4% of the original parameters—specifically transitioning from ChatGPT to Llama 7B (Touvron et al., 2023a). SLIM employs streamlined templates to facilitate this transfer, enabling SLMs to bolster recommender systems by incorporating additional item information. Evaluations across three Amazon review categories—games, food, and home—demonstrate that SLIM not only maintains high recommendation quality but also effectively reduces popularity bias, all while significantly lowering operational costs.
@@
SLM for Long-term Planning. Traditional recommendation systems primarily focus on optimizing users’ immediate responses, often maximizing short-term benefits but neglecting long-term engagement, leading to issues like confining users within an echo chamber of preferred information and filter bubbles (Wang et al., 2022; Gao et al., 2023). To address this, it is crucial to integrate planning abilities into recommendations to consider both immediate and long-term outcomes. Language models, with their world knowledge and reasoning abilities, are anticipated to offer powerful planning capabilities. BiLLP (Shi et al., 2024) introduces a hierarchical learning approach with macro-learning and micro-learning phases. Macro-learning involves a Planner and a Reflector, both implemented as SLM instances such as Llama-2-7B (Touvron et al., 2023b). The Planner uses high-level experiences to create long-term plans, while the Reflector updates the Planner by reflecting on past actions. Micro-learning employs an SLM-based Actor-Critic mechanism for personalized planning, where the Actor converts plans into actions and the Critic evaluates the actions for long-term benefits. Similar to SLMs for cold-start problems, the use of SLMs for long-term planning is also underexplored and merits further investigation.
@@
SLM as Content Encoder. Language models, particularly when deep, provide an effective starting point for fine-tuning on downstream tasks. In news recommendation systems, the representational capability of a model significantly impacts performance. Consequently, many news recommender systems now employ language models fine-tuned on specific datasets as text encoders. For example, Wu et al. (2021) conducts pioneering work using a pre-trained language model to enhance large-scale news recommender systems by substituting traditional news encoders with a BERT model (Devlin et al., 2019). While models such as BERT (Devlin et al., 2019) may struggle to capture essential content when pre-trained on limited data, more versatile LMs such as Llama-2-7B (Touvron et al., 2023b) enhance text descriptions and demonstrate improvements in recommendation performance through advanced embedding techniques. ONCE (Liu et al., 2024a) propose using Llama-2-7B (Touvron et al., 2023b) as an encoder to overcome the limitations of BERT in content-based recommendations. Additionally, the study explores the synergistic use of LLMs such as ChatGPT in recommendation systems, finding that SLMs optimized with LoRA (Hu et al., 2021) outperform the recommendation results of systems assisted by generic LLMs such as ChatGPT.
@@ 4.1.4. SLM Applications in Web Search
Web search systems, involving retrieval and ranking, face challenges due to the diverse web documents and search queries. Traditional keyword-matching methods often fall short because of phrasing variations and the long-tail distribution of queries and content, complicating accurate semantic inference. Effective integration of retrieval and ranking models is also crucial. Language models, serving as content encoders, help overcome semantic challenges through their deep language understanding from pre-training (Chu et al., 2022; Dong et al., 2023; Wang et al., 2024g). Joint training of retrieval and ranking models addresses integration, with SLMs ranking retrieved documents and acting as re-rankers. Additionally, SLMs serve as rewriters in scenarios requiring enhanced query understanding. Thus, in web search, SLMs fulfill three roles: content encoder (Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020), ranker (Chu et al., 2022; Nogueira and Cho, 2019), and rewriter (Ma et al., 2023d), as depicted in Figure 13. Next, we give details.
@@
SLM as a Content Encoder. Text embeddings are vector representations of natural language that encode semantic information, widely used in retrieval. SLM-based embedding retrieval, also known as dense retrieval, leverages deep language understanding from pre-training to effectively address semantic challenges. H-ERNIE (Chu et al., 2022) features a hierarchical model structure that encodes both the query and the document at various levels of granularity, such as character, word, and phrase. This design addresses the ambiguous model understanding in web searches, exemplified by a search query for ”red panda” returning results related to ”panda.” The model includes an aggregation module that collects information from finer-grained layers to construct the coarser-grained layers, enhancing the specificity and relevance of search results. Implicit Interaction ($I^{3}$) (Dong et al., 2023) uses BERT (Devlin et al., 2019) as a content encoder, generating implicit pseudo-queries from passages to enable high online efficiency with offline caching of passage vectors. Similarly, Zou et al. (2022) employs BERT and ERNIE as encoders, adapting them for integration into retrieval systems through techniques like fine-tuning for doc-query similarity tasks. To improve online service responsiveness, they suggest using knowledge distillation to create lightweight service ranking models that preserve effectiveness while meeting latency demands.
However, ERNIE and BERT-style models overlook advancements in SLMs like context length extension. Thus, Wang et al. (2024g) employs the improved Mistral-7B (Jiang et al., 2023) and fine-tunes it on synthetic data across 93 languages generated by GPT-4 (Achiam et al., 2023) to enhance embeddings. These enhanced embeddings show competitive performance on the BEIR (Thakur et al., 2021) and MTEB (Muennighoff et al., 2023) benchmarks, improving by a notable margin (+2%). They also effectively handle personalized passkey retrieval for inputs up to 32k tokens by modifying the rotation base of position embeddings, extending the context well beyond the usual 512-token limit. Peng et al. (2023b) employs LLaMa-7B (Touvron et al., 2023a) and Vicuna-7B (Chiang et al., 2023) as semantic encoders for embedding retrieval, demonstrating improved performance through soft prompt tuning. CoCondenser (Gao and Callan, 2022) addresses sensitivity to noisy data and large batch requirements during dense retriever training. Using the Condenser architecture with Transformer blocks, the model condenses information into dense vectors effectively. CoCondenser applies an unsupervised contrastive loss, minimizing data manipulation and batch size needs. Tests on MS-MARCO (Bajaj et al., 2016), Natural Questions (Kwiatkowski et al., 2019), and Trivia QA (Joshi et al., 2017) demonstrate its robustness and efficacy.
@@
SLM as a Ranker. The reranking task, the following retrieval, improves the order of multiple candidates to enhance retrieval quality because rerankers are more accurate than embedding retrievers. InPars (Inquisitive Parrots for Search) (Bonifacio et al., 2022) introduces a retrieval framework that employs the T5 base 220M (Raffel et al., 2020) as a re-ranker to enhance the BM25 retriever (Robertson et al., 2009). Initially, BM25 selects 1,000 candidates, which are then re-ranked using a fine-tuned T5 model, called monoT5. This T5 model is adapted as a binary classifier to predict the relevance of a document to a query. The training data, generated by GPT-3 (Brown et al., 2020), creates queries for specific documents and selects negative examples randomly. Experimental results demonstrate that this specifically fine-tuned SLM-enhanced retriever significantly outperforms the version utilizing GPT-3 (Brown et al., 2020). For instance, performance on the TREC-DL 2020 dataset (Craswell et al., 2021) shows that retrieval with monoT5-220M achieves a 0.3599 MAP score, surpassing that of GPT-3 (175B), which scores 0.3163.
@@
SLM as a Rewriter. The query to the retriever, often in the form of several keywords, may expose a knowledge gap between the actual query and the knowledge required for effective querying, limiting retrieval performance. For instance, in the retrieval-augmented generation, this gap places a burden on prompt engineering to the downstream LLMs. To address this, the ”rewrite-retrieve-read” framework (Ma et al., 2023d) employs T5-large (Raffel et al., 2020) as a rewriter to bridge the gap between queries and necessary knowledge by rewriting inputs. A specialized, trainable language model, known as the ”rewriter,” is used for this task. It is trained via reinforcement learning, with downstream LLM performance as a reward, to better adapt queries for downstream tasks. Experimental results show that this trainable SLM rewriter exceeds the performance of general LLM rewrites. For example, on HotpotQA, the rewrite-retrieve-read framework achieves a 45.97 F1 score, surpassing the generic LLM’s 43.85 F1 score.

### v2

@@ 4.1.1. SLM Applications in Question-Answering
    Retrieval-Augmented Generation (RAG): Incorporating External Knowledge for Domain-Specific QA. RAG addresses OOD questions by integrating external knowledge during inference, allowing models to access information beyond their pre-trained parameters. By retrieving relevant documents in real time, RAG enables small language models to provide accurate answers on specialized topics. In the telecommunications domain, Gichamba et al. (2024) use ColBERT as a dense retrieval system to fetch documents from technical datasets. By encoding queries and documents separately, ColBERT computes relevance scores, helping small models like Phi-2 and Falcon-7B retrieve precise technical information to answer complex telecom-related queries. Rationale Ranking (Hartill et al., 2023a) addresses answering unseen questions using smaller language models by integrating external explanatory contexts from retrieval systems with reasoning rationales from LLMs. This method involves ranking both the retrieved explanatory contexts and LLM-generated rationales using a scoring module, which then combines them to form a cohesive context. Consequently, this integrated approach enhances the SLMs’ performance on unseen questions.
@@
    Self-Adaptive Techniques: Enhancing Model Adaptability with Self-Generated Pseudo Labels. Fine-tuning, while effective in adapting domain knowledge, can be impractical in realistic scenarios where labeled datasets are scarce. To overcome this, self-adaptive techniques employ self-generated pseudo labels to activate specific aspects of the target tasks, thereby enhancing model adaptability (Veksler, 2023; Shu et al., 2022). Test-time Self-Adaptive Small LMs (T-SAS) (Jeong et al., 2023) first stochastically generates multiple answers for an unlabeled question. The most plausible answer is then selected via majority voting to enhance pseudo-label accuracy, serving as a pseudo-label for training during test-time.
Comparison between LLMs and SLMs for QA. When comparing LLMs such as GPT-4 (Achiam et al., 2023) or BLOOM-175B (Le Scao et al., 2023) with fine-tuned SLMs in QA tasks, the benefits of SLMs are clear. LLMs, while versatile across multiple domains due to extensive pre-training, are computationally demanding, making them less ideal for resource-limited settings. SLMs, however, when fine-tuned for specific domains, often match or exceed the performance of larger models within those specialties. The trade-off is between large-scale models’ generalization and small-scale model’s specialization: LLMs handle diverse domains but may need additional techniques such as knowledge injection for domain-specific queries. In contrast, domain-specific SLMs, though less flexible, provide higher accuracy and more relevant responses, making them ideal for edge deployments where computational resources are scarce but domain precision is crucial.
@@ 4.1.2. SLM Applications in Coding
The adoption of SLMs for coding offers an alternative to LLMs due to their lower computational needs and potential for domain-specific tuning. Despite LLMs’ proficiency in code generation and programming support, SLMs are advantageous for their faster inference, reduced operational costs, and suitability for real-time environments where rapid responses are crucial. Representative works are discussed next. The Phi series (Javaheripi et al., 2023; Abdin et al., 2024; Li et al., 2023b) showcase SLMs’ evolution in coding tasks. For instance, Phi-1 (Gunasekar et al., 2023), a Transformer with 1.3B parameters, specializes in basic Python coding and achieves notable scores in benchmarks such as HumanEval (Gunasekar et al., 2023), which includes 164 programming problems. Subsequent models, Phi-1.5 and Phi-2, have enhanced these capabilities, while Phi-3 demonstrated SLMs’ potential to rival larger models (Abdin et al., 2024). The latest model, Phi-3.5-mini, with 3.8B parameters, excels in long context tasks using advanced fine-tuning and optimization techniques, performing comparably to larger models such as Llama-3.1-8B-instruct (Dubey et al., 2024) and surpassing smaller ones like Gemma-2 (Team et al., 2024b).
@@
Comparison between SLMs and LLMs on Coding. Table 5 provides a comparative analysis of SLMs and LLMs on coding benchmarks HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). Insights include: (i) Small SLMs (1.3B - 3.8B Parameters) like Phi-3.5-mini (Trufinescu, 2024) achieve high scores, demonstrating the efficacy of small models. Mid-sized SLMs (6.7B - 9B Parameters), such as DeepSeek-Coder 6.7B (Guo et al., 2024b) and Llama 3.1 8B (Dubey et al., 2024), show improved performance, indicating that larger model sizes and enhanced training contribute to better accuracy. Large models (33B and above) like Llama 3.1 405B (Dubey et al., 2024), GPT-4o (OpenAI, 2024b), and Claude 3.5 Sonnet (Anthropic, 2024) excel, supporting the idea that bigger models generalize better across diverse coding tasks; (ii) There’s a notable trade-off between computational efficiency and performance, with larger models requiring more resources, impacting their practical deployment in constrained environments; (iii) Specialized training and fine-tuning, as used in models like DeepSeek-Coder (Guo et al., 2024b), are crucial for excelling in coding tasks, though such models may not handle complex requests as effectively, highlighting the versatility of general SLMs for broader applications.
@@ 4.1.3. SLM Applications in Recommender Systems
Recommender systems are essential in various online services, helping to manage information overload and meet users’ personal needs. SLMs enhance recommendation systems by (1) addressing the cold start problem; (2) reducing popularity bias; (3) improving long-term planning; (4) serving as personalized recommenders; and (5) acting as content encoders. These applications show the versatility and effectiveness of SLMs in boosting performance and personalization in recommendation. Next, we introduce the details.
@@
SLM for System Cold Start Problem. Traditional recommendation systems, which utilize historical user-item interactions such as clicks, purchases, and ratings to learn representations and match items to users, fail in scenarios lacking any user-item interactions, known as the cold-start recommendation problem, often occurring in start-up businesses (Rashid et al., 2008). Although LLMs address this with in-context learning, their slow and costly inference restricts real-time use. Thus, PromptRec (Wu et al., 2024d) explores using SLMs as in-context recommenders for recommendation system cold-start problems. However, SLMs often struggle without emergent context-learning abilities. To overcome this, SLMs are enhanced by pre-training on relevant corpora, using a improved C4 corpus subset (Raffel et al., 2020), and by developing training prompts for different domains, enhancing cold-start performance. Results show that enhanced SLMs like BERT-mini (Devlin et al., 2019), with 11.3M parameters, achieve BERT-large’s performance in cold-start scenarios, with only 17% of BERT-large’s inference time. Similarly, many studies have addressed the cold-start problem by leveraging BERT (Zhang et al., 2019; Noorian, 2024; Zhuang and Kim, 2021; Heidari et al., 2022). For example, ADLRS (Heidari et al., 2022) employs BERT to convert web-crawled item profiles into vectors that highlight key aspects, aiding recommender systems in acquiring essential initial information.
SLM for Mitigating Popularity Bias. Popularity bias in recommender systems, marked by discrepancies between item popularity in training datasets and the real world, often stems from using closed-loop datasets with limited information. Recent LLMs leverage their broad open-world knowledge to better reason about user-item interactions (Lin et al., 2024b; Liu et al., 2024a), reducing this bias by providing recommenders with more extensive item details. Using the chain-of-thought (CoT) prompting, LLMs decompose complex tasks into intermediate reasoning steps, enhancing understanding of user behavior and interests. However, LLMs’ high resource demands limit their practical use. To overcome this, the Step-by-step Knowledge Distillation Framework for Recommendation (SLIM) (Wang et al., 2024g) distills LLM reasoning capabilities into SLMs, keeping just 4% of the original parameters, transitioning from ChatGPT to Llama 7B (Touvron et al., 2023a). SLIM uses detailed LLM templates to extract reasoning steps and streamlined templates for fine-tuning, enabling SLMs to improve recommender systems by better reasoning on richer item information.
@@
SLM for Long-term Planning. Traditional recommender systems focus on optimizing immediate user responses, often maximizing short-term gains but overlooking long-term engagement. This can trap users in echo chambers and filter bubbles (Wang et al., 2022b; Gao et al., 2023). To tackle this, integrating planning capabilities into recommendations to balance immediate and long-term outcomes is vital. LMs, with their extensive knowledge and reasoning abilities, are expected to enhance planning capabilities. BiLLP (Shi et al., 2024) adopts a hierarchical learning approach with macro and micro-learning phases. In macro-learning, a Planner and a Reflector, both as SLM instances like Llama-2-7B (Touvron et al., 2023b), operate; the Planner forms long-term plans using high-level experiences, while the Reflector updates plans based on past actions. Micro-learning uses an SLM-based Actor-Critic mechanism for personalized planning, with the Actor implementing plans and the Critic assessing actions for long-term benefits. The use of SLMs for long-term planning, similar to their use in cold-start scenarios, remains underexplored and merits further research.
@@
SLM as a Content Encoder. Language models, particularly when deep, provide an effective starting point for fine-tuning on downstream tasks. In news recommendation systems, the representational capability of a model significantly impacts performance. Consequently, many news recommender systems now employ language models fine-tuned on specific datasets as text encoders. For example, Wu et al. (2021) conducts pioneering work using a pre-trained language model to enhance large-scale news recommender systems by substituting traditional news encoders with a BERT model (Devlin et al., 2019). However, BERT may struggle to capture content as it is pre-trained on limited data. Therefore, ONCE (Liu et al., 2024a) propose using Llama-2-7B (Touvron et al., 2023b) as an encoder to overcome the limitations of BERT in content-based recommendations. Additionally, the study explores the synergistic use of LLMs in recommendation systems, finding that SLMs optimized with LoRA (Hu et al., 2021) outperform the recommendation results of systems assisted by generic LLMs such as ChatGPT.
@@ 4.1.4. SLM Applications in Web Search
Web search systems, involving retrieval and ranking, face challenges due to the diverse web documents and search queries. Traditional keyword-matching methods often fall short because of phrasing variations and the long-tail distribution of queries and content, complicating accurate semantic inference. Effective integration of retrieval and ranking models is also crucial. Language models, serving as content encoders, help overcome semantic challenges through their language understanding from pre-training (Chu et al., 2022; Dong et al., 2023; Wang et al., 2024h). Joint training of retrieval and ranking models addresses integration, with SLMs ranking retrieved documents and acting as re-rankers. Additionally, SLMs serve as rewriters in scenarios requiring enhanced query understanding. Thus, in web search, SLMs fulfill three roles: (1) content encoder, (2) ranker, and (3) rewriter, as depicted in Figure 15. Next, we give details.
@@
SLM as a Content Encoder. Text embeddings are vector representations that encode semantic information, widely used in retrieval; SLM-based dense retrieval utilizes pre-trained deep language understanding to effectively tackle semantic challenges. H-ERNIE (Chu et al., 2022) employs a hierarchical model that encodes queries and documents at multiple granularity—character, word, and phrase—to improve specificity and relevance in web search results by aggregating finer details into coarser layers, addressing issues like ambiguous queries. Implicit Interaction ($I^{3}$) (Dong et al., 2023) uses BERT (Devlin et al., 2019) as a content encoder, generating implicit pseudo-queries from passages to enable high online efficiency with offline caching of passage vectors. However, ERNIE and BERT-style models overlook advancements in SLMs such as context length extension (Roziere et al., 2023). Thus, Peng et al. (2023c) employs LLaMa-7B (Touvron et al., 2023a) and Vicuna-7B (Chiang et al., 2023) as semantic encoders for embedding retrieval, demonstrating improved performance through soft prompt tuning. CoCondenser (Gao and Callan, 2022) addresses sensitivity to noisy data and large batch requirements during dense retriever training. Using the Condenser architecture with Transformer blocks, the model condenses information into dense vectors effectively.
@@
SLM as a Ranker. The reranking task improves the order of multiple candidates to enhance retrieval quality because rerankers are more accurate than embedding retrievers. InPars (Inquisitive Parrots for Search) (Bonifacio et al., 2022) employs the T5 base 220M (Raffel et al., 2020) as a re-ranker to enhance the BM25 retriever (Robertson et al., 2009). Initially, BM25 selects 1K candidates, re-ranked by a fine-tuned T5 model (monoT5) adapted as a binary classifier to assess document-query relevance. Training data, generated by GPT-3 (Brown et al., 2020), formulates queries and selects random negative examples. Experiments show the monoT5-enhanced retriever significantly outperforms GPT-3; for example, it achieves a 0.3599 MAP score on the TREC-DL 2020 dataset (Craswell et al., 2021), surpassing GPT-3’s 0.3163.
@@
SLM as a Rewriter. Queries to the retriever, typically just a few keywords, may reveal a knowledge gap between the query and the knowledge needed for effective retrieval, thus limiting performance. To address this, the “rewrite-retrieve-read” framework (Ma et al., 2023d) uses T5-large (Raffel et al., 2020) to bridge the knowledge gap in queries by rewriting them for more effective retrieval. This rewriter, trained via reinforcement learning with downstream LLM performance as a reward, outperforms general LLM rewrites. For example, it achieves a 45.97 F1 score on HotpotQA, surpassing the generic LLM’s 43.85 F1 score.